# Long-Tailed Continual Learning For Visual Food Recognition

### v1

@@ III Long-tailed Food Datasets For Diabetes
Figure 2: The distribution of VFN-INSULIN and VFN-T2D shown in descending order based on number of training samples.
Figure 3: The overview of our proposed framework. The red arrow shows the training procedure where at each incremental learning phase, the input training data contains the new classes images and the exemplars for all seen classes. We leverage the balanced softmax loss [45] $\mathcal{L}_{bs}$ based on the distribution of input training data to learn class-balance representation. We also employed a fixed teacher model obtained from the last incremental step to perform feature-based knowledge distillation $\mathcal{L}_{fkd}$ to address catastrophic forgetting. The blue arrows denote the steps after the training process where we select a small fixed number of $\mathcal{M}$ exemplars per class and store in the memory buffer. Finally, before the next learning phase, we perform the CAM-based exemplar augmentation to construct a balanced exemplar set and combine with the new classes data for the next training phase.
@@ IV Method
In this work, we introduce a novel end-to-end long-tailed continual learning framework for visual food recognition. The overview of our method is shown in Figure 3. To address catastrophic forgetting, we leverage the teacher model learned from the last incremental step and perform feature-based knowledge distillation with an additional prediction head to enable the efficient knowledge transfer. The exemplar set selected based a novel CAM-based data augmentation for tail classes. Finally, we replace the cross-entropy with the balanced softmax loss [45] based on the current training data distribution to learn class-balance visual representation. In this section, we first introduce the preliminaries for continual learning in long-tailed distribution in Section IV-A and then illustrate the detail of each proposed components in Section IV-BIV-C and IV-D, respectively.
@@ IV-A Preliminaries
We focus on continual learning in class-incremental settings where the objective is to learn new classes incrementally and perform classification on all classes seen so far during the inference phase. Specifically, the continual learning in the class-incremental scenario can be formulated as applying an initial model $h_{0}$ to learn a sequence of $N$ tasks denoted as $\mathcal{T}=\{\mathcal{T}^{1},...,\mathcal{T}^{N}\}$ where each task $\mathcal{T}^{i}$ contains $C_{i}$ non-overlapped new classes, which is also known as the incremental step size. During the learning phase of each new task, only the training data $D_{i}=\{\textbf{x}_{i}^{j},y_{i}^{j}\}$ of the current task is available where $\textbf{x}_{i}^{j}$ and $y_{i}^{j}$ denote the $j$-th input image and label, respectively. After each incremental learning step, the updated model $h_{i}$ needs to classify $C_{1:i}$ classes encountered so far. The major challenge of continual learning is catastrophic forgetting [15] where the updated model $h_{i}$ after learning the task $\mathcal{T}^{i}$ forgets the knowledge of previous tasks $\{\mathcal{T}^{1},...,\mathcal{T}^{i-1}\}$, resulting in significant performance degradation to classify $C_{1:i-1}$. In the conventional setup, the training data $D_{i}$ for each task $\mathcal{T}^{i}$ is balanced distributed, containing $|D_{i}|/C_{i}$ samples per class. However, this assumption simplifies the real world complexities especially for food recognition where data is usually long-tailed distributed and exhibits imbalance among food classes. Formally, the training data $D_{i}$ for each task in long-tailed continual learning is a class-imbalanced distributed with each class containing $(0,|D_{i}|)$ training samples. The entire training data $D$ for all the $N$ tasks $\mathcal{T}$ exhibits the long-tail distribution.
@@ IV-A1 Knowledge distillation
Finally, the knowledge distillation is integrated with cross-entropy during the training process by using a hyper-parameter $\alpha$ to learn new classes as well as maintaining the learned knowledge.
@@ IV-A2 Exemplar replay
As one of the most commonly used strategies to address catastrophic forgetting, the exemplar replay based methods [60, 61, 57] assume the availability of a reasonable memory budget to select a small fixed number of data as exemplars for each seen class and store them in memory buffer (also known as exemplar set). Specifically, after learning each task $\mathcal{T}^{i}$, the lower layers of updated model $h_{i}$ is used to extract feature embeddings for the new classes training data $D_{i}=\{\textbf{x}_{i}^{j},y_{i}^{j}\}$. The Herding algorithm [59] is widely applied to select the most representative data for each new class based on euclidean distance between feature embedding and the class mean vector. Therefore, given a memory budget of $\mathcal{M}$ data per class (also known as memory capacity), a subset of $E_{i}\subseteq D_{i}$ is selected with $|E_{i}|=\mathcal{M}\times C_{i}$ and stored in the memory buffer. Finally, at the beginning of the next new task $\mathcal{T}^{i+1}$, all the exemplars in memory buffer are combined with the new classes training data to construct $E_{i}+D_{i+1}$ for continual learning. In this work, we use Herding as the exemplar selection algorithm while others latest work [42, 67] could also be applied.
@@ IV-B Feature-based Knowledge Distillation
Despite the effectiveness of knowledge distillation in conventional continual learning setup as described in Section IV-A1, it is difficult to apply when new classes exhibit long-tailed distribution since the output logits of the teacher model could be heavily biased towards instance-rich classes due to the severe class-imbalance problem [62, 63]. Directly applying knowledge distillation as in (2) on biased output logits may even harm the overall performance [19]. Therefore, we explore the feature-based knowledge distillation in this work for more efficient knowledge transfer in the long-tailed continual learning scenario. As discussed in [20], one of the challenges when applying feature-based distillation is called the feature space misalignment where the representation of student and teacher models could mismatch in terms of both magnitude and direction. This problem is also relevant in the long-tailed continual learning scenario as the new knowledge is acquired sequentially, which requires the evolution of the continual learning model in the feature space to accommodate the addition of new classes. In this work, we address this challenge by introducing a simple yet effective method as shown in Figure 4. Specifically, instead of directly mimicking the feature from teacher model, we apply an additional predictor $g$ on the head of the continual learning model to map the current representation space to the past in the teacher model. Given an image x, the predictor $g$ takes the feature representation from the current model (i.e. student model) $h_{i}(\textbf{x})$ as input and outputs the mapped feature $g(h_{i}(\textbf{x}))$. Then we distill the knowledge from the teacher model $h_{i-1}$ by
@@
where <,><,>< , > measures the cosine similarity. By applying the predictor, we provide the student model with more freedom to accommodate the previous learned representation into the current feature space, enabling more efficient knowledge distillation in long-tailed continual learning. The predictor $g$ is be removed after each incremental learning phase. Note that although we apply cosine embedding loss for knowledge distillation, our method can be integrated with other loss functions such as the Mean Squared Error (MSE) loss.
@@ IV-C CAM-based Exemplar Augmentation
Existing exemplar replay based methods assume each class should contain at least $\mathcal{M}$ images given $\mathcal{M}$ as the memory budget in Section IV-A2. However, most classes in long-tailed distribution may contain only a few training samples $n<\mathcal{M}$, which imposes two new challenges including (i) inefficiency of knowledge replay due to the insufficient of training samples and (ii) intensification of the class-imbalance issue if we directly combine the stored exemplars with training data from new class due to the imbalanced nature of memory buffer. Therefore, we propose a novel data augmentation method in this work to construct a balanced exemplar set by augmenting the tail classes images to address both aforementioned issues. The overview of proposed data augmentation technique is illustrated in Figure 5. To address the issue of losing semantic information when performing data augmentation [21, 14] as described in Section II-B, we propose to use a class activation map (CAM) [68] to identify the most important region from instance-rare classes images and then preserve the semantic information by performing CutMix [22] to cut and paste the identified region into the images with rich context that are selected based on visual similarity. Specifically, we construct class-balanced memory buffer before each new task $\mathcal{T}^{i+1}$ by augmenting stored images for food classes $C_{t}$ with less than $\mathcal{M}$ exemplars through CutMix in conjunction with images selected from food classes $C_{h}$ containing $\mathcal{M}$ exemplars. Given an input image $\textbf{x}_{t}\in C_{t}$, we first select the most visually similar candidate $\textbf{x}_{h}\in C_{h}$ by comparing the cosine similarity with $h_{i}$ as feature extractor where $\textbf{x}_{h}=\underset{\textbf{x}_{k}\in C_{h}}{\text{argmax}}<h_{i}(\textbf{x}_{t}),h_{i}(\textbf{x}_{k})>$. The lower half of Figure 5 illustrates the procedure to identify the the region to cut and paste into $\textbf{x}_{h}$. Formally, given $\textbf{x}_{t}\in\mathbb{R}^{c\times h\times w}$, the class-activation map $M(\textbf{x}_{t})\in\mathbb{R}^{h\times w}$ is calculated by
@@
where $v_{y_{\textbf{x}_{t}}}\in\mathbb{R}^{d}$ refers to the weight vector in classifier of the current model corresponding to the seen class $y_{\textbf{x}_{t}}\in C_{1:i}$. The value of CAM ranges from $[0,1]$ and a higher value indicates the more discriminative class-specific region. Therefore, we apply a random threshold $\sigma\in(0,1)$ to select the region $M(\textbf{x}_{t})^{T}\in\mathbb{R}^{h\times w}$ where
@@
without losing semantic information of the input image. Finally, we perform CutMix to generate a synthetic image $\tilde{\textbf{x}}_{t}$ by
@@
where $A_{r}$ and $A$ denote the area of the replaced region and the total area of $\textbf{x}_{h}$. $y_{h}$ and $y_{t}$ refer to the original class label for $\textbf{x}_{h}$ and $\textbf{x}_{t}$, respectively. Note that the exemplar augmentation is performed at the beginning of each new task and the augmented images are not stored in the memory buffer. Note that the Grad-CAM[69], which can be regarded as the generalization of CAM [68], could also be applied in our method.
Figure 5: The overview of proposed CAM-based data augmentation technique. The green arrow describes the selection of the most visually similar candidate image and the red arrow illustrates the steps to obtain the most important region of input image to perform CutMix [22].
@@ IV-D Integrated Loss
While the exemplar augmentation mitigates the class-imbalance issue by constructing a balanced memory buffer, the number of available training data between new classes and the stored classes may still vary a lot during the training phase due to the limited memory budget. Existing work [17] addresses this problem by decoupling the training process into two stages to first learn a feature extractor and then fine-tune the classifier using class-balanced sampler. In this work, we propose to use Balanced Softmax (BS) [45] by extending it into a long-tailed continual learning scenario without requiring a decoupled training process. Specifically, during the training phase of the new task $\mathcal{T}^{i}$, a distribution vector $v_{d}\in\mathbb{R}^{C_{1:i}}$ is generated by counting the number of training data of each food class for input images in the current task. Recall $L\in\mathbb{R}^{C_{1:i}}$ is the output logits from current model $h_{i}$, the distribution vector is then used as the prior information when calculating the loss as shown in (10)
@@
where $\bar{v}_{d}=v_{d}/sum(v_{d})$ is the normalized distribution vector and $\Phi()$ denotes the $Softmax$ function. Therefore, the larger value in the distribution vector achieves smaller gradients when we compute the cross-entropy using the adjusted logits $v_{t}+L$ and vice versa, which address the class-imbalance issue and enables the end-to-end training pipeline.
@@
TABLE II: Results on Food101-LT, VFN-LT, VFN-INSULIN and VFN-T2D by comparing with existing continual learning methods in terms of last step accuracy ($A_{L}$) and avergae accuracy $A_{M}$. Best results are marked in bold.
Datasets Food101-LT  VFN-LT    VFN-INSULIN   VFN-T2D
Number of tasks $N=5$ $N=10$ $N=20$ $N=7$ $N=7$ $N=7$
Accuracy (%) $A_{L}$ $A_{M}$ $A_{L}$ $A_{M}$ $A_{L}$ $A_{M}$ $A_{L}$ $A_{M}$ $\quad A_{L}\quad$ $A_{M}$ $A_{L}$ $A_{M}$
LwF [55] 8.62 10.02 5.86 10.10 0.83 3.58 1.02 4.85 4.69 12.35 3.80 7.10
EWC [54] 4.29 5.05 3.70 7.23 0.83 3.96 1.58 6.31 2.97 12.05 1.50 4.47
iCaRL [60] 11.48 12.42 12.46 13.42 11.04 12.77 11.84 18.76 1.91 4.49 10.64 12.36
EEIL [61] 12.88 12.68 10.57 14.63 6.98 12.66 16.87 20.77 14.77 17.36 13.02 12.95
LwM [70] 10.62 10.82 7.22 12.27 2.45 6.82 7.41 12.32 6.24 15.47 4.24 8.13
IL2M [19] 12.55 11.45 10.97 14.21 6.81 12.16 14.87 18.68 14.19 15.24 10.71 11.93
BiC [62] 14.94 16.72 12.39 16.22 10.38 14.36 15.40 20.89 13.53 19.37 15.32 20.70
LUCIR [57] 13.87 16.94 10.17 15.75 6.74 9.32 16.31 23.37 15.35 20.68 13.42 16.38
PODNet [58] 12.04 12.22 10.46 12.07 8.99 12.41 14.75 18.21 14.18 15.06 12.78 15.55
EEIL-2stage [17] 14.16 14.96 13.29 16.38 9.76 14.71 17.61 22.98 13.64 18.60 12.86 14.47
LUCIR-2stage [17] 16.34 18.90 13.03 18.75 10.85 16.53 17.33 24.26 14.49 22.49 12.82 17.85
PODNet-2stage [17] 13.94 17.89 11.12 17.21 10.28 16.28 17.14 25.58 16.04 22.28 14.24 19.71
Ours 17.54 21.83 13.57 19.25 11.78 17.43 20.43 29.33 18.49 27.51 18.54 28.17

### v2

@@ III-B Long-tailed Food Datasets For Different Populations
Figure 1: The distribution of VFN186-LT, VFN186-INSULIN, and VFN186-T2D shown in descending order based on the number of training samples.
The process of generating the long-tailed datasets follows  [12] to reduce the number of training samples for food classes in the original VFN186 based on the matched consumption frequency. Overall, VFN186-LT contains 5,185 training images across 186 classes, VFN186-INSULIN contains 4,179 training images, and VFN186-T2D contains 4,403 training images, with a maximum of 324 and a minimum of 1 image per class. The imbalance ratio $\rho$, defined as the maximum over the minimum number of training samples, equals $324$ for three datasets. The food type Yeast bread is highly consumed among represented adults and dominates the consumption frequency in all groups, while frequencies of other food types vary between them. Figure 1 shows the distribution of food types in VFN186-LT, VFN186-INSULIN, and VFN186-T2D, ranked by the number of training samples per class.
Figure 2: The overview of our proposed framework. The red arrows show the training process with new class images and exemplars from previous classes. The blue arrows denote the steps after the training process where we construct a balanced exemplar set and store them in the memory buffer.
@@ IV Method
In this work, we introduce a novel end-to-end long-tailed continual learning framework for visual food recognition. The overview of our method is shown in Figure 2. To address catastrophic forgetting, we leverage the teacher model learned from the last incremental step and perform feature-based knowledge distillation with an additional prediction head to enable efficient knowledge transfer. The exemplar set was selected based on a novel CAM-based data augmentation for tail classes. Finally, we replace the cross-entropy with the balanced softmax loss [50] based on the current training data distribution to learn class-balance visual representation. In this section, we first introduce the preliminaries for continual learning in the long-tailed distribution in Section IV-A and then illustrate the detail of each proposed component in Section IV-BIV-C and IV-D, respectively.
@@ IV-A Preliminaries
We focus on continual learning in class-incremental settings where the objective is to learn new classes incrementally and perform classification on all classes seen so far during the inference phase. Specifically, the continual learning in the class-incremental scenario can be formulated as applying an initial model $h_{0}$ to learn a sequence of $N$ tasks denoted as $\mathcal{T}=\{\mathcal{T}^{1},...,\mathcal{T}^{N}\}$ where each task $\mathcal{T}^{i}$ contains $C_{i}$ non-overlapped new classes, which is also known as the incremental step size. During the learning phase of each new task, only the training data $D_{i}=\{\textbf{x}_{i}^{j},y_{i}^{j}\}$ of the current task is available where $\textbf{x}_{i}^{j}$ and $y_{i}^{j}$ denote the $j$-th input image and label, respectively. After each incremental learning step, the updated model $h_{i}$ needs to classify $C_{1:i}$ classes encountered so far. The major challenge of continual learning is catastrophic forgetting [14] where the updated model $h_{i}$ after learning the task $\mathcal{T}^{i}$ forgets the knowledge of previous tasks $\{\mathcal{T}^{1},...,\mathcal{T}^{i-1}\}$, resulting in significant performance degradation to classify $C_{1:i-1}$. In the conventional setup, the training data $D_{i}$ for each task $\mathcal{T}^{i}$ is evenly distributed, containing $|D_{i}|/C_{i}$ samples per class. However, this assumption simplifies the real-world complexities, especially for food recognition where data is usually long-tailed distributed and exhibits imbalance among food classes. Formally, the training data $D_{i}$ for each task in long-tailed continual learning is a class-imbalanced distribution with each class containing $(0,|D_{i}|)$ training samples. The entire training data $D$ for all the $N$ tasks $\mathcal{T}$ exhibits the long-tail distribution.
@@ IV-A1 Knowledge distillation
Finally, the knowledge distillation is integrated with cross-entropy during the training process by using a hyper-parameter $\alpha$ to learn new classes and maintain the learned knowledge.
@@ IV-A2 Exemplar replay
As one of the most commonly used strategies to address catastrophic forgetting, the exemplar replay-based methods [64, 66, 61] assume the availability of a reasonable memory budget to select a small fixed number of data as exemplars for each seen class and store them in memory buffer (also known as exemplar set). Specifically, after learning each task $\mathcal{T}^{i}$, the lower layers of updated model $h_{i}$ are used to extract feature embeddings for the new classes training data $D_{i}=\{\textbf{x}_{i}^{j},y_{i}^{j}\}$. The Herding algorithm [63] is widely applied to select the most representative data for each new class based on the Euclidean distance between feature embedding and the class mean vector. Therefore, given a memory budget of $\mathcal{M}$ data per class (also known as memory capacity), a subset of $E_{i}\subseteq D_{i}$ is selected with $|E_{i}|=\mathcal{M}\times C_{i}$ and stored in the memory buffer. Finally, at the beginning of the next new task $\mathcal{T}^{i+1}$, all the exemplars in the memory buffer are combined with the new classes training data to construct $E_{i}+D_{i+1}$ for continual learning. In this work, we use Herding as the exemplar selection algorithm while other latest work [48] could also be applied.
@@ IV-B Feature-based Knowledge Distillation
Despite the effectiveness of knowledge distillation in conventional continual learning setup as described in Section IV-A1, it is difficult to apply it to long-tailed distributions since the output logits of the teacher model can be heavily biased towards instance-rich classes [67]. Directly applying knowledge distillation as in (1) on biased output logits may even harm the overall performance [25]. Therefore, we explore feature-based knowledge distillation for better knowledge transfer in long-tailed continual learning. However, a key challenge is feature space misalignment of the challenges when applying feature-based distillation, where the representation of student and teacher models could mismatch in terms of both magnitude and direction [26]. This problem is also relevant in continual learning as the model evolves to incorporate new classes. To solve this, we introduce a simple yet effective method as shown in Figure 3. Specifically, instead of directly mimicking the feature from the teacher model, we apply an additional predictor $g$ on the head of the continual learning model to map the current representation space to the past in the teacher model. $g$ is a single-layer perceptron that performs domain mapping while preserving consistent dimensions. Specifically, it maps from $\mathbb{R}^{d\times 1}$ to $\mathbb{R}^{d\times 1}$, followed by a ReLU activation function. The dimensional consistency is crucial to ensure that the student model, with the added $g$, still outputs image features of the same size, i.e., $d\times 1$. Given an image x, the predictor $g$ takes the feature representation from the current model (i.e. student model) $h_{i}(\textbf{x})$ as input and outputs the mapped feature $g(h_{i}(\textbf{x}))$. Then we distill the knowledge from the teacher model $h_{i-1}$ by
@@
where <,><,>< , > measures the cosine similarity. By applying the predictor, we give the student model more freedom to accommodate the previously learned representation into the current feature space, enabling more efficient knowledge distillation in long-tailed continual learning. The predictor $g$ is removed after each incremental learning phase. Note that although we apply cosine embedding loss for knowledge distillation, our method can be integrated with other loss functions such as the Mean Squared Error (MSE) loss.
@@ IV-C CAM-based Exemplar Augmentation
Existing exemplar replay-based methods assume each class should contain at least $\mathcal{M}$ images given $\mathcal{M}$ as the memory budget in Section IV-A2. However, most classes in long-tailed distribution may contain only a few training samples $n<\mathcal{M}$, which imposes two new challenges including (i) inefficiency of knowledge replay due to the insufficient training samples and (ii) intensification of the class-imbalance issue if we directly combine the stored exemplars with training data from new class due to the imbalanced nature of memory buffer. Therefore, we propose a novel data augmentation method in this work to construct a balanced exemplar set by augmenting the tail class images to address both aforementioned issues. The overview of the proposed data augmentation technique is illustrated in Figure 4. To address the issue of losing semantic information when performing data augmentation [27, 12] as described in Section II-B, we propose to use a class activation map (CAM) [76] to identify the most important region from instance-rare classes images and then preserve the semantic information by performing CutMix [28] to cut and paste the identified region into the images with rich context that are selected based on visual similarity. Specifically, we construct a class-balanced memory buffer before each new task $\mathcal{T}^{i+1}$ by augmenting stored images for food classes $C_{t}$ with less than $\mathcal{M}$ exemplars through CutMix in conjunction with images selected from food classes $C_{h}$ containing $\mathcal{M}$ exemplars. Given an input image $\textbf{x}_{t}\in C_{t}$, we first select the most visually similar candidate $\textbf{x}_{h}\in C_{h}$ by comparing the cosine similarity with $h_{i}$ as feature extractor where $\textbf{x}_{h}=\underset{\textbf{x}_{k}\in C_{h}}{\text{argmax}}<h_{i}(\textbf{x}_{t}),h_{i}(\textbf{x}_{k})>$. The lower half of Figure 4 illustrates the procedure to identify the region to cut and paste into $\textbf{x}_{h}$. Formally, given $\textbf{x}_{t}\in\mathbb{R}^{c\times h\times w}$, the class-activation map $M(\textbf{x}_{t})\in\mathbb{R}^{h\times w}$ is calculated by
@@
where $v_{y_{\textbf{x}_{t}}}\in\mathbb{R}^{d}$ refers to the weight vector in the classifier of the current model corresponding to the seen class $y_{\textbf{x}_{t}}\in C_{1:i}$. The value of CAM ranges from $[0,1]$ and a higher value indicates the more discriminative class-specific region. Therefore, we apply a random threshold $\sigma\in(0,1)$ to select the region $M(\textbf{x}_{t})^{T}\in\mathbb{R}^{h\times w}$ where
@@
without losing the semantic information of the input image. Finally, we apply CutMix to generate a synthetic image $\tilde{\textbf{x}}_{t}$ by
@@
where $A_{r}$ and $A$ denote the area of the replaced region and the total area of $\textbf{x}_{h}$, and $y_{h}$ and $y_{t}$ are the original class labels of $\textbf{x}_{h}$ and $\textbf{x}_{t}$. The exemplar augmentation is performed at the beginning of each new task and the augmented images are not stored in the memory buffer. Note that the Grad-CAM[77], which can be regarded as the generalization of CAM [76], could also be applied in our method.
Figure 4: The overview of proposed CAM-based data augmentation technique. The green arrow describes the selection of the most visually similar candidate image and the red arrow illustrates the steps to obtain the most important region of the input image to perform CutMix [28].
@@ IV-D Integrated Loss
While the exemplar augmentation mitigates the class-imbalance issue by constructing a balanced memory buffer, the number of available training data between new classes and the stored classes may still vary a lot during the training phase due to the limited memory budget. Existing work [22] addresses this problem by decoupling the training process into two stages to first learn a feature extractor and then fine-tune the classifier using a class-balanced sampler. In this work, we propose to use Balanced Softmax (BS) [50] by extending it into a long-tailed continual learning scenario without requiring a decoupled training process. Specifically, during the training phase of the new task $\mathcal{T}^{i}$, a distribution vector $v_{d}\in\mathbb{R}^{C_{1:i}}$ is generated by counting the number of training data of each food class for input images in the current task. Recall $L\in\mathbb{R}^{C_{1:i}}$ is the output logits from current model $h_{i}$, the distribution vector is then used as the prior information when calculating the loss as shown in (9)
@@
where $\bar{v}_{d}=v_{d}/sum(v_{d})$ is the normalized distribution vector and $\Phi()$ denotes the $Softmax$ function. Therefore, the larger value in the distribution vector achieves smaller gradients when we compute the cross-entropy using the adjusted logits $v_{t}+L$ and vice versa. This addresses the class-imbalance issue and enables the end-to-end training pipeline.
@@
[…]
Datasets Food101-LT VFN-LT VFN186-LT VFN186-INSULIN VFN186-T2D ImageNetSubset-LT
Number of tasks $N=5$ $N=10$ $N=20$ $N=7$ $N=9$ $N=9$ $N=9$ $N=10$ $N=20$
LwF [60] 10.02 5.86 0.83 4.85 11.60 10.99 11.12 23.55 20.16
EWC [59] 5.05 3.70 0.83 6.31 11.09 10.40 4.12 19.73 16.49
iCaRL [64] 12.42 12.46 11.04 18.76 8.76 7.23 7.56 33.75 29.71
LwM [78] 10.82 7.22 2.45 12.32 11.04 10.37 10.88 30.24 26.63
IL2M [25] 11.45 10.97 6.81 18.68 11.23 10.52 11.30 31.70 25.20
BiC [67] 16.72 12.39 10.38 20.89 9.09 9.27 11.08 33.31 30.86
EEIL-2stage [22] 14.96 13.29 9.76 22.98 12.86 11.74 12.69 36.84 30.39
LUCIR-2stage [22] 18.90 13.03 10.85 24.26 15.80 13.64 14.88 39.87 34.79
PODNet-2stage [22] 17.89 11.12 10.28 25.58 16.00 13.77 15.11 34.79 31.71
MAFDRC [79] 19.04 16.20 13.63 22.48 18.59 17.03 18.54 40.01 34.48
DGR [80] 23.08 20.35 16.43 26.11 19.58 17.66 17.90 45.12 40.79
Ours 27.52 25.12 21.72 27.53 22.21 20.61 21.23 44.68 40.31